how much pure acid must be mixed into 100ml of 20% acid to create a solution of 60% acid?
Answer by mananth(16949)

-------------- percent ---------------- Amount
Acid solution 20 ---------------- 100 ml
Pure acid 100 ---------------- x ml
Mixture 60 ---------------- 100 + x ml

20*100 +100x =60(100 +x)
2000+ 100x=6000+60x
100x-60x=-2000+ 6000
40 x = 4000
/ 40
x = 100 ml Pure acid

m.ananth@hotmail.ca